# Generalization in Federated Learning: A Conditional Mutual Information Framework

### v1

@@ 7 Empirical Verification
In this section, we empirically evaluate the effectiveness of our CMI bounds by comparing them to the expected generalization gap. Specifically, we estimate the fast-rate e-CMI bound for FL, as presented in Theorem 5.1, referred to as the Fast-rate Bound in our experiments. Additionally, due to the challenges associated with estimating MI when dealing with high-dimensional random variables, we compute an evaluated version of the CMI bound from Theorem 4.1. This e-CMI bound, derived in Theorem G.1 (in Appendix), serves as a lower bound for Theorem 4.1 and is denoted as the Square-root Bound in our experiments. Our experimental setup, particularly the construction of supersamples, closely follows previous works (Harutyunyan et al., 2021; Hellström & Durisi, 2022a; Wang & Mao, 2023b, 2024; Dong et al., 2024), and we construct the superclient in a similar manner. For the FL learning process, we implement FedAvg following the federated training protocol of McMahan et al. (2017), with modifications to reduce computational overhead. We conduct image classification experiments on two datasets: MNIST (LeCun et al., 2010) and CIFAR-10 (Krizhevsky, 2009).
Experiment Setup
@@
For both of the classification tasks, clients train locally for five epochs per round before sending their models to the central server, with training spanning $300$ communication rounds (which reduces the commonly used $1000$ rounds to lower computational costs). Additionally, we apply a pathological non-IID data partitioning scheme as in McMahan et al. (2017), and we evaluate prediction error as our performance metric, i.e. we utilize the zero-one loss function to compute generalization error. During training, we use cross-entropy loss as a surrogate to enable optimization with gradient-based methods. Since we pre-define the non-participating clients from the superclient, all participating clients are included in every round of training.
The complete experiment details including CMI estimation can be found in Appendix H.
@@
Results
Figure 1 presents the estimated expected generalization gap alongside our CMI bounds. Notably, both CMI bounds are non-vacuous, meaning they remain below the upper bound of the loss function (i.e., $1$). More importantly, they effectively capture the generalization behavior of the FL learning algorithm and are numerically very close to the expected generalization gap. For MNIST, where the algorithm achieves a small training error, the fast-rate bound from Theorem 5.1 performs significantly better than the square-root CMI bound in Theorem 4.1 in this low empirical risk regime. On CIFAR-10, since we use a simple CNN model and limit the number of communication rounds, the empirical risk remains non-negligible. In this case, we observe that the evaluated version of Theorem 4.1 outperforms Theorem 5.1 in most settings. While the tightness of e-CMI bounds in standard centralized i.i.d. learning has been observed in previous works, our results confirm that CMI bounds remain numerically tight in FL.
@@ Appendix C Input-Output Mutual Information (IOMI) Generalization Bound
To further tighten the generalization bound for $\mathcal{E}(\mathcal{A})$, we adopt techniques introduced by Bu et al. (2020) to derive a tighter bound for $\mathcal{E}{OG}(\mathcal{A})$, along with similar developments from Wang & Mao (2023a) for bounding $\mathcal{E}{PG}(\mathcal{A})$.
@@
$|\mathcal{E}(\mathcal{A})|\leq\frac{1}{K}\sum_{i=1}^{K}\sqrt{2\sigma^{2}\mathrm{D_{KL}}(P_{Z}||\mu_{i})}+\frac{1}{nK}\sum_{i=1}^{K}\sum_{j=1}^{n}\sqrt{2\sigma^{2}I(W;Z_{i,j})},$ (3)
@@
The proof of this theorem directly follows from the techniques in Bu et al. (2020, Proposition 1) and Wang & Mao (2023a, Theorem 4.1).
Compared to the CMI bounds, this bound has its own distinctive advantage: the first term explicitly characterizes how the divergence between client distributions affects the participation gap. This provides a clearer understanding of how heterogeneity in client distributions impacts the generalization performance of FL algorithms.
@@ Appendix F Omitted Proof in Section 6
We apply the loss-difference (LD)-CMI technique from Wang & Mao (2023a) to derive a tighter version of Theorem 4.1.

### v2

@@ 6 Empirical Verification
In this section, we empirically evaluate the effectiveness of our CMI bounds by comparing them to the expected generalization gap. Specifically, we estimate the fast-rate e-CMI bound for FL, as presented in Theorem 4.1, referred to as the Fast-rate Bound in our experiments. Additionally, due to the challenges associated with estimating MI when dealing with high-dimensional random variables, we compute an evaluated version of the CMI bound from Theorem 3.1. This e-CMI bound, derived in Theorem G.1 (in Appendix), serves as a lower bound for Theorem 3.1 and is denoted as the Square-root Bound in our experiments. Our experimental setup, particularly the construction of supersamples, closely follows previous works (Harutyunyan et al., 2021; Hellström & Durisi, 2022a; Wang & Mao, 2023c, 2024b; Dong et al., 2024), and we construct the superclient in a similar manner. For the FL learning process, we implement FedAvg following the federated training protocol of McMahan et al. (2017), with modifications to reduce computational overhead. We conduct image classification experiments on two datasets: MNIST (LeCun et al., 2010) and CIFAR-10 (Krizhevsky, 2009). The details of our experimental setup and results are presented below.
6.1 Experiment Setup
@@
For both of the classification tasks, clients train locally for five epochs per round before sending their models to the central server, with training spanning $300$ communication rounds (which reduces the commonly used $1000$ rounds to lower computational costs). Additionally, we apply a pathological non-IID data partitioning scheme as in McMahan et al. (2017): data are sorted by label, split into $200$ shards of size $300$, and each client is randomly assigned $2$ shards, and we evaluate prediction error as our performance metric, i.e. we utilize the zero-one loss function to compute generalization error. During training, we use cross-entropy loss as a surrogate to enable optimization with gradient-based methods. Since we pre-define the non-participating clients from the superclient, all participating clients are included in every round of training.
The complete experiment details including other hyperparameter settings and CMI estimation can be found in Appendix H.
@@
6.2 Results
Figure 1 presents the estimated expected generalization gap alongside our CMI bounds. Notably, both CMI bounds are non-vacuous, meaning they remain below the upper bound of the loss function (i.e., $1$). More importantly, they effectively capture the generalization behavior of the FL learning algorithm and are numerically very close to the expected generalization gap. For MNIST, where the algorithm achieves a small training error, the fast-rate bound from Theorem 4.1 performs significantly better than the square-root CMI bound in Theorem 3.1 in this low empirical risk regime. On CIFAR-10, since we use a simple CNN model and limit the number of communication rounds, the empirical risk remains non-negligible. In this case, we observe that the evaluated version of Theorem 3.1 outperforms Theorem 4.1 in most settings. While the tightness of e-CMI bounds in standard centralized i.i.d. learning has been observed in previous works (Harutyunyan et al., 2021; Hellström & Durisi, 2022a; Wang & Mao, 2023c; Dong et al., 2024), our results confirm that CMI bounds remain numerically tight in FL.
7 Other Related Literature
Federated Learning Methods
FedAvg (McMahan et al., 2017) is one of the earliest and most widely used FL algorithms. However, its performance degrades significantly in non-i.i.d. settings. To mitigate this issue, FedProx (Li et al., 2020) introduces a regularization term, $||w-w_{i}||$, which helps stabilize updates and reduce divergence among local models. In fact, the squared form of this regularization term naturally arises in the derivation of our Theorem 5.2 (see Appendix F.2), and when the CMI term in Theorem 5.2 is small, this regularization term also tends to be small. Another approach, SCAFFOLD (Karimireddy et al., 2020), tackles data heterogeneity using variance reduction techniques to improve convergence. FedOpt (Reddi et al., 2021) extends adaptive optimization methods to FL, enabling more efficient learning dynamics, while FedNova (Wang et al., 2020) introduces a normalized averaging scheme to mitigate objective inconsistencies while preserving fast training convergence. More recently, leveraging information-theoretic analysis, FedMDMI (Zhang et al., 2024) incorporates a global model-data MI term as a regularizer to improve generalization performance. Beyond these methods, numerous other FL algorithms have been proposed, we refer readers to (Huang et al., 2024; Solans et al., 2024) for recent advances.
Generalization Theory in Federated Learning
Recent research has increasingly focused on understanding generalization in FL. In addition to works in (Chen et al., 2021; Hu et al., 2023; Yagli et al., 2020; Barnes et al., 2022; Gholami & Seferoglu, 2024; Zhang et al., 2024), which we have discussed throughout the paper, several other studies provide new insights. Sefidgaran et al. (2022, 2024) introduce rate-distortion-based generalization bounds for FL, with Sefidgaran et al. (2024) demonstrating that increasing the frequency of communication rounds does not always reduce generalization error and, in some cases, may even degrade generalization performance. Stability-based generalization analysis has also been explored in FL. For instance, Sun et al. (2024) analyze the impact of data heterogeneity, concluding that greater heterogeneity leads to higher generalization error. It is also worth noting that generalization theory in meta-learning is closely related to FL. Particularly, personalized FL can be viewed as a variant of meta-learning (Fallah et al., 2020). In fact, our CMI framework is inspired in part by a recent CMI-based generalization bound for meta-learning proposed in Hellström & Durisi (2022b). However, our results are not directly comparable to theirs due to a key difference in problem setup: their meta-learning framework requires the meta-learner (i.e., the global model $W$) to be further trained on the test tasks (i.e., previously non-participating clients), whereas in FL, the global model is evaluated directly on unseen clients without any additional local fine-tuning. To enable a direct comparison with Hellström & Durisi (2022b), our CMI framework would need to be extended to the personalized FL setting, where local adaptation of the global model is permitted.
@@ 8 Concluding Remarks
Acknowledgements
The authors would like to thank the anonymous AC and reviewers for their careful reading and valuable suggestions. The authors acknowledge the use of GPT-4o to assist in polishing the language of this work. All AI-generated text was carefully reviewed and edited by the authors, who take full responsibility for the final content of the paper.
Impact Statement
The primary focus of our work is on advancing the theoretical foundations of federated learning, aiming to deepen our understanding of this field. As with most theoretical work, it inherently presents challenges in explicitly highlighting potential societal consequences. Meanwhile, the proposed theoretical framework, particularly in privacy constraint based results, may find application in related research domains like trustworthy AI. While the theoretical results in such domains could have positive societal impacts, it’s crucial to acknowledge the potential for misuse. The research community plays a vital role in delivering guidance on responsible and ethical applications to navigate the evolving landscape of technology responsibly.
@@ Appendix C Input-Output Mutual Information (IOMI) Generalization Bound
To further tighten the generalization bound for $\mathcal{E}_{OG}(\mathcal{A})$, we adopt techniques introduced by Bu et al. (2020), along with similar developments from Wang & Mao (2023b) for bounding $\mathcal{E}_{PG}(\mathcal{A})$.
@@
$|\mathcal{E}_{\mathcal{D}}(\mathcal{A})|\leq\frac{1}{K}\sum_{i=1}^{K}\sqrt{2\sigma^{2}\mathrm{D_{KL}}(P_{Z}||\mu_{i})}+\frac{1}{nK}\sum_{i=1}^{K}\sum_{j=1}^{n}\sqrt{2\sigma^{2}I(W;Z_{i,j})},$ (3)
@@
The proof of this theorem follows directly from the techniques in Bu et al. (2020, Proposition 1) and Wang & Mao (2023b, Theorem 4.1).
Compared to the CMI bounds, this bound has its own distinctive advantage: the first term explicitly characterizes how the divergence between client distributions affects the participation gap. This provides a clearer understanding of how heterogeneity in client distributions impacts the generalization performance of FL algorithms. In addition, this result will ultimately lead to a gradient-norm-based regularizer when SGD or SGLD is used (Wang & Mao, 2024a) and hints at a potential feature alignment mechanism in the KL sense for clients.
@@ Appendix F Omitted Proof in Section 5
We apply the loss-difference (LD)-CMI technique from Wang & Mao (2023b) to derive a tighter version of Theorem 3.1.